OpenStudy (anonymous):

Medale For Help - Medal for help :) Solve the system of equations- y=2x2 -3 y=3x-1 A. no solution b.(-1/2,5), (2,-5/2) c.(-1/2,-5/2),(2,5) d(1/2,5/2),(2,5)

OpenStudy (anonymous):

A.?

OpenStudy (whpalmer4):

\[y = 2x^2-3\]\[y = 3x-1\]Are those the equations?

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

Then no, A is not correct. Here's how you can solve it. The solution to the system of equations will be a pair of numbers (x,y) that satisfy both equations. Your equations are both written \(y = \text{<stuff>}\) and the value of \(y\) is the same, so we can just set the other two sides of the equation equal to each other: \[2x^2-3 = 3x-1\] Does that make sense?

OpenStudy (anonymous):

Its c ( -1/2, -5/2 ) ( 2 , 5 )

OpenStudy (whpalmer4):

That's a quadratic equation which you can solve in a number of different ways.

OpenStudy (whpalmer4):

start by moving all the terms to one side: \[2x^2-3 - 3x = 3x - 3x - 1\]\[2x^2-3-3x+1 = 3x-3x-1+1\]\[2x^2-3x-2 = 0\]Probably the easiest thing here is to use the quadratic formula for the solution of \(ax^2+bx+c=0\):\[x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]